# On the Mean Value of a Weighted Composite Arithmetic Function

### v1

@@ Abstract
The primary objective of this paper is to employ methods from analytic number theory to investigate the mean value properties of a composite function involving the Dirichlet divisor function and a generalized minimal power function. Specifically, we study the weighted summatory function where the divisor function is normalized by the number of distinct prime factors. We establish a rigorous asymptotic formula for this sum, detailing the analytic properties of the associated Dirichlet series and the contour integration process.
@@ 1 Introduction
Let $r\geq 2$ be a fixed integer. For any positive integer $n$, we define the generalized minimal power function of order $r$, denoted by $\mathcal{M}_{r}(n)$, as:
@@
$D_{k}(n)=\frac{d(n)}{k^{\omega(n)}}.$ (3)
Since both $d(n)$ and $k^{\omega(n)}$ are multiplicative functions, $D_{k}(n)$ is also multiplicative. Note that for $k=1$, $D_{1}(n)$ reduces to the standard divisor function $d(n)$. The weight $k^{-\omega(n)}$ serves to dampen the contribution of integers with many distinct prime factors.
@@
$\sum_{n\leq X}D_{k}(\mathcal{M}_{r}(n))=\mathcal{C}(r,k)x\ln x+\mathcal{K}x+O\left(x^{\frac{1}{2}+\epsilon}\right),$ (4)
@@ 2 Preliminary Lemmas
Let $f(n)$ be a multiplicative function such that the series $\sum_{n=1}^{\infty}|f(n)|n^{-\sigma}$ converges for $\sigma>1$. Then:
$\sum_{n=1}^{\infty}\frac{f(n)}{n^{s}}=\prod_{p}\left(1+\sum_{a=1}^{\infty}\frac{f(p^{a})}{p^{as}}\right),\quad\Re(s)>1.$ (5)
@@
Let $F(s)=\sum_{n=1}^{\infty}a_{n}n^{-s}$ be a Dirichlet series with abscissa of convergence $\sigma_{c}$. For $X>0$, $X\notin\mathbb{Z}$, $b>\max(0,\sigma_{c})$, and $T>0$:
$\sum_{n\leq X}a_{n}=\frac{1}{2\pi i}\int_{b-iT}^{b+iT}F(s)\frac{X^{s}}{s}ds+O\left(\sum_{n=1}^{\infty}|a_{n}|\left(\frac{X}{n}\right)^{b}\min\left(1,\frac{1}{T|\ln(X/n)|}\right)\right).$ (6)
@@ 3 Proof of the Main Result
$f(s)=\sum_{n=1}^{\infty}\frac{F_{r,k}(n)}{n^{s}}=\sum_{n=1}^{\infty}\frac{d(\mathcal{M}_{r}(n))}{k^{\omega(n)}n^{s}}.$ (8)
Based on the analysis of the multiplicative properties of $F_{r,k}(n)$, this series admits an Euler product expansion. It can be meromorphically continued to the half-plane $\text{Re}(s)>1/2$, possessing a pole of order 2 at $s=1$. This structure is analogous to the square of the Riemann zeta-function, $\zeta(s)^{2}$, modified by convergent Euler products.
@@
$R(x,T)\ll\sum_{n=1}^{\infty}\left|F_{r,k}(n)\right|\left(\frac{x}{n}\right)^{b}\min\left(1,\frac{1}{T\left|\ln\frac{x}{n}\right|}\right)+\frac{|F_{r,k}(x)|}{T}.$ (10)
@@
To extract the asymptotic behavior, we shift the line of integration to the left, into the region where $\text{Re}(s)$ is smaller, while capturing the singularities of the integrand. Consider the rectangular contour $\mathcal{C}$ defined by the vertices biTb-iT, $b+iT$, $\sigma_{0}+iT$, and $\sigma_{0}-iT$, where $\sigma_{0}=\frac{1}{2}+\epsilon$. This choice is dictated by the region of absolute convergence of the analytic part of $f(s)$ and the known bounds of the Riemann zeta-function. According to the Residue Theorem, the integral over the original vertical segment is equal to the sum of the residues of the poles located inside the contour, plus the integrals along the other three segments of the rectangle
$\int_{b-iT}^{b+iT}=2\pi i\sum\text{Res}+\int_{\text{top}}+\int_{\text{left}}+\int_{\text{bottom}}.$ (11)
The function $f(s)\frac{x^{s}}{s}$ has a pole of order 2 at $s=1$ inside this contour (since $b>1$ and $\sigma_{0}<1$).
This is the crucial step that determines the main terms of the asymptotic formula ($x\ln x$ and $x$). In the neighborhood of $s=1$, the function $f(s)$ behaves like $\zeta(s)^{2}$ multiplied by an analytic function. Since $\zeta(s)$ has a simple pole at $s=1$ with residue 1, $f(s)$ has a pole of order 2. We write the Laurent expansion
$f(s)=\frac{A}{(s-1)^{2}}+\frac{B}{s-1}+O(1),\quad\text{as }s\to 1.$ (12)
The constants $A$ and $B$ appearing in the Laurent expansion of $f(s)$ at $s=1$ are explicitly computable from the Euler product structure of the Dirichlet series. Specifically, $A$ corresponds to the leading constant $\mathcal{C}(r,k)$ in our Theorem.
@@
$H(s)=\frac{\zeta(ks)}{\zeta(2s)}\prod_{p}\left(1-\frac{1}{k(p^{ks}+p^{(k-1)s})}\right).$ (14)
@@
$A=\frac{6\zeta(k)}{\pi^{2}}\prod_{p}\left(1-\frac{1}{k(p^{k}+p^{k-1})}\right).$ (16)
@@
$\zeta(s)=\frac{1}{s-1}+\gamma+O(s-1),$ (17)
@@
$\zeta(s)^{2}=\frac{1}{(s-1)^{2}}+\frac{2\gamma}{s-1}+O(1).$ (18)
@@
$B=H^{\prime}(1)+2\gamma H(1)=H^{\prime}(1)+2\gamma A.$ (19)
@@
$\frac{H^{\prime}(1)}{H(1)}=\frac{\zeta^{\prime}(k)}{\zeta(k)}\cdot k-2\frac{\zeta^{\prime}(2)}{\zeta(2)}+\sum_{p}\frac{d}{ds}\left[\ln\left(1-\frac{1}{k(p^{ks}+p^{(k-1)s})}\right)\right]_{s=1}.$ (20)
Each term in this expression is explicitly computable: $\zeta(k)$ and $\zeta^{\prime}(k)$ are known values for integer $k\geq 2$, $\zeta(2)=\frac{\pi^{2}}{6}$, $\zeta^{\prime}(2)=\frac{\pi^{2}}{6}(\gamma+\ln(2\pi)-12\ln A)$ where $A$ is the Glaisher-Kinkelin constant, and the sum over primes converges rapidly.
@@
$\text{Res}_{s=1}\left(f(s)\frac{x^{s}}{s}\right)=Ax\ln x+(B-A)x=\mathcal{C}(r,k)x\ln x+\mathcal{K}x,$ (21)
where:
$\mathcal{C}(r,k)=A=\frac{6\zeta(k)}{\pi^{2}}\prod_{p}\left(1-\frac{1}{k(p^{k}+p^{k-1})}\right),$ (22)
and:
$\mathcal{K}=B-A=H^{\prime}(1)+(2\gamma-1)A.$ (23)
@@
$\left|\int_{\sigma_{0}}^{b}f(\sigma\pm iT)\frac{x^{\sigma\pm iT}}{\sigma\pm iT}\,d\sigma\right|\ll\frac{x^{b}}{T}\int_{\sigma_{0}}^{b}|f(\sigma\pm iT)|\,d\sigma.$ (24)
@@
$\int_{\sigma_{0}-iT}^{\sigma_{0}+iT}f(s)\frac{x^{s}}{s}\,ds.$ (25)
@@
$\int_{-T}^{T}\left|f\left(\frac{1}{2}+\epsilon+it\right)\right|\frac{x^{\frac{1}{2}+\epsilon}}{|t|}\,dt\ll x^{\frac{1}{2}+\epsilon}\cdot T^{\epsilon}\ll x^{\frac{1}{2}+\epsilon}.$ (26)
@@
$\frac{1}{2\pi i}\int_{b-iT}^{b+iT}f(s)\frac{x^{s}}{s}\,ds=\text{Res}_{s=1}+O\left(x^{\frac{1}{2}+\epsilon}\right).$ (27)
Substituting this back into the original Perron’s formula in (9), we rigorously establish the theorem
$\sum_{n\leq x}F_{r,k}(n)=\mathcal{C}(r,k)x\ln x+\mathcal{K}x+O\left(x^{\frac{1}{2}+\epsilon}\right).$ (28)

### v2

@@ Abstract
The main objective of this paper is to use methods from analytic number theory to study the mean value properties of a composite function involving the Dirichlet divisor function and a generalised minimal power function. More specifically, we study the weighted sum function where the divisor function is normalised by the number of distinct prime factors. We establish a rigorous asymptotic formula for this sum, detailing the analytical properties of the associated Dirichlet series and the contour integration process. The main result provides an explicit expansion with a dominant term depending on $X\ln X$ and $X$ and an error term of order $O(X^{\frac{1}{2}+\epsilon})$.
@@ 1 Introduction
Let $r\geq 2$ be a fixed integer. For any positive integer $n$, we define the generalized minimal power function of order $r$, denoted by $\mathcal{M}_{r}(n)$, as
@@
$D_{k}(n)=\frac{d(n)}{k^{\omega(n)}}.$
Since both $d(n)$ and $k^{\omega(n)}$ are multiplicative functions, $D_{k}(n)$ is also multiplicative. Note that for $k=2$, the weighted divisor function $D_{2}(n)=\frac{d(n)}{2^{\omega(n)}}$ is closely related to the arithmetic function $\frac{d(n)}{d^{*}(n)}$ studied in recent literature [2], where $d^{*}(n)$ denotes the number of unitary divisors of $n$. The weight $k^{-\omega(n)}$ serves to dampen the contribution of integers with many distinct prime factors.
@@
$\sum_{n\leq x}D_{k}(\mathcal{M}_{r}(n))=\mathcal{C}(r,k)x\ln x+\mathcal{K}x+O\left(x^{\frac{1}{2}+\epsilon}\right),$
@@ 2 Preliminary Lemmas
Let $f(n)$ be a multiplicative function such that the series $\sum_{n=1}^{\infty}|f(n)|n^{-\sigma}$ converges for $\sigma>1$. Then
$\sum_{n=1}^{\infty}\frac{f(n)}{n^{s}}=\prod_{p}\left(1+\sum_{a=1}^{\infty}\frac{f(p^{a})}{p^{as}}\right),\quad\Re(s)>1.$
@@
Let $F(s)=\sum_{n=1}^{\infty}a_{n}n^{-s}$ be a Dirichlet series with abscissa of convergence $\sigma_{c}$. For $X>0$, $X\notin\mathbb{Z}$, $b>\max(0,\sigma_{c})$, and $T>0$
$\sum_{n\leq X}a_{n}=\frac{1}{2\pi i}\int_{b-iT}^{b+iT}F(s)\frac{X^{s}}{s}ds+O\left(\sum_{n=1}^{\infty}|a_{n}|\left(\frac{X}{n}\right)^{b}\min\left(1,\frac{1}{T|\ln(X/n)|}\right)\right).$
@@ 3 Proof of the Main Result
$f(s)=\sum_{n=1}^{\infty}\frac{F_{r,k}(n)}{n^{s}}=\sum_{n=1}^{\infty}\frac{d(\mathcal{M}_{r}(n))}{k^{\omega(n)}n^{s}}.$
Based on the analysis of the multiplicative properties of $F_{r,k}(n)$, this series admits an Euler product expansion. For $\text{Re}(s)>1$, we have
$f(s)=\prod_{p}\left(1+\sum_{a=1}^{\infty}\frac{d(\mathcal{M}_{r}(p^{a}))}{k^{\omega(p^{a})}p^{as}}\right).$
For any prime power $p^{a}$ with $a\geq 1$, we have $\omega(p^{a})=1$ and $d(\mathcal{M}_{r}(p^{a}))=\lceil a/r\rceil+1$. Thus, the local factor at prime $p$ becomes
$1+\frac{1}{k}\sum_{a=1}^{\infty}(\lceil a/r\rceil+1)(p^{-s})^{a}.$
To evaluate the infinite sum explicitly, we apply Lemma 6. Setting $z=p^{-s}$, where $|z|<1$ for $\text{Re}(s)>0$, Lemma 6 yields
$\sum_{a=1}^{\infty}(\lceil a/r\rceil+1)z^{a}=\frac{z(2-z^{r})}{(1-z)(1-z^{r})}.$
Substituting this back into the Euler product, the local factor becomes
$1+\frac{1}{k}\left(\frac{p^{-s}(2-p^{-rs})}{(1-p^{-s})(1-p^{-rs})}\right).$
This explicit form allows us to isolate the dominant poles. Through some manipulation and comparison with Euler’s products of the Riemann zeta function, we can factor the general series as follows
$f(s)=\zeta(s)^{2}\cdot H(s),$
where $H(s)$ is analytic and non-zero in the region $\text{Re}(s)\geq\frac{1}{2}+\epsilon$. Specifically, $H(s)$ contains the remaining convergent factors
$H(s)=\frac{\zeta(rs)}{\zeta(2s)}\prod_{p}\left(1-\frac{1}{k(p^{rs}+p^{(r-1)s})}\right).$
This factorization confirms that $f(s)$ possesses a pole of order 2 at $s=1$, arising from the $\zeta(s)^{2}$ term.
@@
$R(x,T)\ll\sum_{n=1}^{\infty}\left|F_{r,k}(n)\right|\left(\frac{x}{n}\right)^{b}\min\left(1,\frac{1}{T\left|\ln\frac{x}{n}\right|}\right)+\frac{|F_{r,k}(x)|}{T}.$
@@
To extract the asymptotic behavior, we shift the line of integration to the left, into the region where $\text{Re}(s)$ is smaller, while capturing the singularities of the integrand. Consider the rectangular contour $\mathcal{C}$ defined by the vertices biTb-iT, $b+iT$, $\sigma_{0}+iT$, and $\sigma_{0}-iT$, traversed in the counter-clockwise direction, where $\sigma_{0}=\frac{1}{2}+\epsilon$. This choice is dictated by the region of absolute convergence of the analytic part of $f(s)$ and the known bounds of the Riemann zeta-function.
Let $\Gamma_{1}$ denote the original segment from biTb-iT to $b+iT$, and let $\Gamma_{2},\Gamma_{3},\Gamma_{4}$ be the upper, left, and lower segments, respectively, closing the contour. According to Cauchy’s residue theorem, the integral over the closed contour $\mathcal{C}=\Gamma_{1}+\Gamma_{2}+\Gamma_{3}+\Gamma_{4}$ satisfies:
$\oint_{\mathcal{C}}f(s)\frac{x^{s}}{s}\,ds=2\pi i\sum_{\rho\in\text{int}(\mathcal{C})}\text{Res}_{s=\rho}\left(f(s)\frac{x^{s}}{s}\right).$
Consequently, the integral over the original vertical segment $\Gamma_{1}$ can be expressed as
$\int_{b-iT}^{b+iT}f(s)\frac{x^{s}}{s}\,ds=2\pi i\sum_{\rho\in\text{int}(\mathcal{C})}\text{Res}_{s=\rho}\left(f(s)\frac{x^{s}}{s}\right)-\left(\int_{\Gamma_{2}}+\int_{\Gamma_{3}}+\int_{\Gamma_{4}}\right)f(s)\frac{x^{s}}{s}\,ds.$
The function $f(s)\frac{x^{s}}{s}$ has a unique singularity inside this contour at $s=1$. According to the factorization $f(s)=\zeta(s)^{2}H(s)$, where $H(s)$ is analytic at $s=1$, the integrand has a pole of order 2 at $s=1$ (since $b>1$ and $\sigma_{0}<1$). Thus, the sum of residues reduces to the single term $\text{Res}_{s=1}\left(f(s)\frac{x^{s}}{s}\right)$.
This is the crucial step that determines the main terms of the asymptotic formula ($x\ln x$ and $x$). In the neighborhood of $s=1$, the function $f(s)$ behaves like $\zeta(s)^{2}$ multiplied by an analytic function. Since $\zeta(s)$ has a simple pole at $s=1$ with residue 1, $f(s)$ has a pole of order 2. We now want to develop the function $f(s)\frac{x^{s}}{s}$.
First, we expand the term $\frac{x^{s}}{s}$ in a Taylor series in the neighborhood of $s=1$. Let $s=1+h$, where $h\to 0$. We have
$x^{s}=x^{1+h}=x\cdot e^{h\ln x}=x\left(1+h\ln x+O(h^{2})\right),$
and
$\frac{1}{s}=\frac{1}{1+h}=1-h+O(h^{2}).$
Multiplying these expansions yields
$\frac{x^{s}}{s}=x\left(1+h\ln x+O(h^{2})\right)\left(1-h+O(h^{2})\right)\\ =x\left(1+h(\ln x-1)+O(h^{2})\right).$
Substituting h=s1h=s-1, we obtain the expansion in terms of $s$
$\frac{x^{s}}{s}=x\left(1+(s-1)(\ln x-1)\right)+O\left((s-1)^{2}\right).$
From the analytic properties of the Dirichlet series established earlier, $f(s)$ admits the following Laurent expansion near $s=1$
$f(s)=\frac{A}{(s-1)^{2}}+\frac{B}{s-1}+O(1),\quad\text{as }s\to 1.$
where $A$ and $B$ are constants determined by the Euler product of $f(s)$.
The residue of $\left(f(s)\frac{x^{s}}{s}\right)$ at $s=1$ is the coefficient of the $(s-1)^{-1}$ term in the product of the two expansions above. Multiplying the series
$f(s)\frac{x^{s}}{s}=\left(\frac{A}{(s-1)^{2}}+\frac{B}{s-1}+O(1)\right)\cdot\left(x+x(\ln x-1)(s-1)+O\left((s-1)^{2}\right)\right)$
$\frac{Ax}{(s-1)^{2}}+\frac{Ax(\ln x-1)}{s-1}+\frac{Bx}{s-1}+O(1).$
Grouping the terms with $(s-1)^{-1}$, we find
$\text{Res}_{s=1}\left(f(s)\frac{x^{s}}{s}\right)=Ax(\ln x-1)+Bx=Ax\ln x+(B-A)x.$
@@
$A=\frac{6\zeta(k)}{\pi^{2}}\prod_{p}\left(1-\frac{1}{k(p^{k}+p^{k-1})}\right).$
@@
$\zeta(s)=\frac{1}{s-1}+\gamma+O(s-1),$
@@
$\zeta(s)^{2}=\frac{1}{(s-1)^{2}}+\frac{2\gamma}{s-1}+O(1).$
@@
$B=H^{\prime}(1)+2\gamma H(1)=H^{\prime}(1)+2\gamma A.$
@@
$\frac{H^{\prime}(1)}{H(1)}=\frac{\zeta^{\prime}(k)}{\zeta(k)}\cdot k-2\frac{\zeta^{\prime}(2)}{\zeta(2)}+\sum_{p}\frac{d}{ds}\left[\ln\left(1-\frac{1}{k(p^{ks}+p^{(k-1)s})}\right)\right]_{s=1}.$
Each term in this expression is explicitly computable: $\zeta(k)$ and $\zeta^{\prime}(k)$ are known values for integer $k\geq 2$, $\zeta(2)=\frac{\pi^{2}}{6}$, $\zeta^{\prime}(2)=\frac{\pi^{2}}{6}(\gamma+\ln(2\pi)-12\ln\lambda)$ where $\lambda\approx 1.2824$ is the Glaisher-Kinkelin constant, and the sum over primes converges rapidly.
@@
$\text{Res}_{s=1}\left(f(s)\frac{x^{s}}{s}\right)=Ax\ln x+(B-A)x=\mathcal{C}(r,k)x\ln x+\mathcal{K}x,$
where
$\mathcal{C}(r,k)=A=\frac{6\zeta(k)}{\pi^{2}}\prod_{p}\left(1-\frac{1}{k(p^{k}+p^{k-1})}\right),$
and
$\mathcal{K}=B-A=H^{\prime}(1)+(2\gamma-1)A.$
@@
$\left|\int_{\sigma_{0}}^{b}f(\sigma\pm iT)\frac{x^{\sigma\pm iT}}{\sigma\pm iT}\,d\sigma\right|\ll\frac{x^{b}}{T}\int_{\sigma_{0}}^{b}|f(\sigma\pm iT)|\,d\sigma.$
@@
$\int_{\sigma_{0}-iT}^{\sigma_{0}+iT}f(s)\frac{x^{s}}{s}\,ds.$
@@
$\int_{-T}^{T}\left|f\left(\frac{1}{2}+\epsilon+it\right)\right|\frac{x^{\frac{1}{2}+\epsilon}}{|t|}\,dt\ll x^{\frac{1}{2}+\epsilon}\cdot T^{\epsilon}\ll x^{\frac{1}{2}+\epsilon}.$
@@
$\frac{1}{2\pi i}\int_{b-iT}^{b+iT}f(s)\frac{x^{s}}{s}\,ds=\text{Res}_{s=1}+O\left(x^{\frac{1}{2}+\epsilon}\right).$
Substituting this back into the original Perron’s formula, we rigorously establish the result of the theorem
$\sum_{n\leq x}F_{r,k}(n)=\mathcal{C}(r,k)x\ln x+\mathcal{K}x+O\left(x^{\frac{1}{2}+\epsilon}\right).$